OpenStudy (anonymous):

HELP, not strong in math: f'x=ln(1-ln(x))

OpenStudy (shubhamsrg):

Need to find f(x) ? Integration ?

OpenStudy (anonymous):

Sorry, new to "open study" The function is fx=ln(1-lnx) I need to find the f' then solve for zero to find critical points.

OpenStudy (anonymous):

maxima minima u mean

OpenStudy (anonymous):

yes

OpenStudy (shubhamsrg):

you are having problems in applying the chain rule ?

OpenStudy (anonymous):

Well, my best guess so far has been 1/(1-lnx) *1. And I think that is correct, but not sure how to solve for zero from here?

OpenStudy (shubhamsrg):

1/(1-lnx) part is correct. WIth that you need to multiple the derivative of (1-lnx) too. So there is a mistake in your derivative. Please re-check.

OpenStudy (anonymous):

Applying the chain rule to the inside (1-ln x) would be one, yes? The 1 would be zero, then would not the "x" in the remaining ln x be "1" or 1/x?

OpenStudy (shubhamsrg):

D(1-lnx) = -1/x :|

OpenStudy (anonymous):

So the derivative of fx=ln(1-ln x) would be: 1/(1-lnx) * -1/x ?

OpenStudy (shubhamsrg):

Yes.

OpenStudy (anonymous):

Thank you, another silly question then, how would the fractions line up?

OpenStudy (shubhamsrg):

Didn't get your query ?

OpenStudy (anonymous):

 trying to insert a drawing of the Equation.